OpenStudy (anonymous):

Use identities (no calculators) to find the exact value for (sin 15)(cos 45) + (cos 15)(sin 45).

OpenStudy (anonymous):

This is the 'addition angle formula \[sin(a+b)=sin(a)cos(b)+cos(a)sin(b)\]

OpenStudy (anonymous):

here \[a=15, b=45\] so the left hand side is \[sin(15+45)=sin(60)\]

OpenStudy (anonymous):

\[sin(60)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Thanks! What if they are angles such as Use identities (no calculators) to find the exact value for (sin 38)(sin 112) -( cos 38)(cos 112) ???